2(x+4) = 4(5-3x): 2(x+4) = 4(5-3x)
Answer by checkley77(3654)

2(x+4)=4(5-3x)
2x+8=20-12x
2x+12x=20-8
14x=12
x=12/14
x=6/7 answer.
proof:
2(6/7+4)=4(5-3*6/7)
2(6+4*7)/7=4(5-18/7)
2(6+28)/7=4(5*7-18)/7
2(34/7)=4(35-18)/7
68/7=4(17/7)
68/7=68/7